If 11, -24, 32, and -9 are the coordinates of the points M, N, P, and Q on a number line, then which of these segments is the longest?
MN, MP, NP, PQ, NQ

Answer by checkley77(12844)

11+24=35 for MN.
32-11=21 for MP.
24+32=56 for NP. is the longest line.
32+9=41 for PQ.
24-9=15 for NQ.